OpenStudy (anonymous):

Purple paint is made with 16:3:1 ratio of white paint: blue paint :red paint. How much white paint in gallons is needed in order to make one gallon of purple paint? Express your answer as a common fraction.

OpenStudy (anonymous):

Will give a medal and fan who ever helps

jimthompson5910 (jim_thompson5910):

"Purple paint is made with 16:3:1 ratio of white paint: blue paint :red paint" so we have 16x gallons of white paint 3x gallons of blue paint 1x gallons of red paint they all mix to get purple. The variable x is some unknown number at this point

jimthompson5910 (jim_thompson5910):

they mix to make 1 full gallon, so 16x + 3x + 1x = 1 solve that for x and then evaluate 16x to get your answer

OpenStudy (anonymous):

Thank you so much for helping

jimthompson5910 (jim_thompson5910):

no problem

OpenStudy (anonymous):

Would that be 18 then?

jimthompson5910 (jim_thompson5910):

16x + 3x + 1x = 1 20x = 1 x = ???

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

sorry it's just im a slow learner

jimthompson5910 (jim_thompson5910):

thats fine

jimthompson5910 (jim_thompson5910):

if 20x = 1, then what is x?

OpenStudy (anonymous):

um 20x=1 so -20 -20 x=-19

jimthompson5910 (jim_thompson5910):

you don't subtract 20 from both sides

jimthompson5910 (jim_thompson5910):

20x means 20 times x so you have to do the opposite of multiplication

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

I'm sorryy

OpenStudy (anonymous):

so then 20x=1 /20 x=1

jimthompson5910 (jim_thompson5910):

you divide BOTH sides by 20 though

jimthompson5910 (jim_thompson5910):

not just that left side

OpenStudy (anonymous):

oh Sorryyy T-T

OpenStudy (anonymous):

20x=1 /20 /20 x=.05

jimthompson5910 (jim_thompson5910):

correct so x = 1/20 or x = 0.05

jimthompson5910 (jim_thompson5910):

you use this x value to compute 16x to get your final answer

OpenStudy (anonymous):

Ohh ok

jimthompson5910 (jim_thompson5910):

it says "Express your answer as a common fraction"

OpenStudy (anonymous):

So then you would get 0.8? which is 80/100

jimthompson5910 (jim_thompson5910):

so you will have to stick with x = 1/20

OpenStudy (anonymous):

oh

jimthompson5910 (jim_thompson5910):

reduce all fractions

jimthompson5910 (jim_thompson5910):

0.8 = 80/100 = ???

OpenStudy (anonymous):

4/5

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

you'll need 4/5 of a gallon of white paint

OpenStudy (anonymous):

Thank you so much for helping me and dealing with my slowness

jimthompson5910 (jim_thompson5910):

that's fine and you're welcome